OpenStudy (anonymous):

find the slope of the line perpendicular to the line whose equation is given by 4y-9x=7. then find the equation of line passing through (3,4)

OpenStudy (anonymous):

1. solve 4y-9x=7 for y 2. take the number beside x (coefficient) and writes its negative reciprocal for example negative reciprocal of 1/2 is -2/1 because its flipped upside down and the sign is changes 3. Then with you new slope (Which is the slope of your new line) put it in point slope for which is y-(the y in the point) = Slope (x - (x in the point)) 4. Solve for y and TA DA! you have the equation of your new line

OpenStudy (anonymous):

@TAKEBACKMATH it looks cool verbal explanation but could you please explain mathematically :)

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@ParthKohli @Mikael

OpenStudy (anonymous):

@everyone please

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

the slope of the given line can be found by comparing it with y=mx+c. 4y-9x=7 so 4y=9x+7 y= 9/4 x + 7/4 slope = 9/4 product of slopes of perpendicular lines = -1 so slope of required line will be -4/9. so u have y= (-4/9)x+c put, (x,y)=(3,4) here and find c. can u?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

4=(-4/9)*3+C ...on solving i get C= 16? right?

OpenStudy (anonymous):

??@hartnn

OpenStudy (anonymous):

thanks :)

hartnn (hartnn):

no!

OpenStudy (anonymous):

what??

hartnn (hartnn):

its c=16/3 :P check again.

hartnn (hartnn):

when u multiply 3 ,it gets multiplied with c also!

hartnn (hartnn):

so finally, your equation will be y=(-4/9)x+16/3 or 9y+4x=48.

OpenStudy (anonymous):

\[4=\frac{ -4 }{ 9 }*3+C..... 4=\frac{ -4 }{ 3 }+C\] oh am so stupid....sorry @hartnn thanks a lot

hartnn (hartnn):

welcome again :)